A mixture initially contains A, B, and C in the following concentrations: [A] = 0.700 M , [B] = 0.850 M , and [C] = 0.400 M . The following reaction occurs and equilibrium is established: A+2B⇌C At equilibrium, [A] = 0.560 M and [C] = 0.540 M . Calculate the value of the equilibrium constant, Kc.
Answer #1

Since the concentration of C increases and A decreases, the reaction has procceded in forward direction.

Change in concentration of B = 2 * change in concentration in a = 2 * (0.7 - 0.56) = 0.280 M

[b] at equilibrium = initial - change in [B] = 0.850 - 0.280 = 0.570 M

Writing the equation for Kc = [c] / ( [A] * [B]2) = 0.540 / ( 0.560 * 0.5702) = 2.968 M-2